# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With Affine Transformation Contrastive Learning

### v3

@@ 5.2.3 Scene Classification
The comparison results are shown in Tab.IX. Although RingMo-Aerial does not perform as well as specialized satellite models like SkySense in nadir-view classification, it still demonstrates strong performance, reflecting its robust generalization capability.
5.2.4 Change Detection
TABLE X: Change Detection: Comp. on LEVIR-CD and CDD
Methods LEVIR-CD CDD
F1 IoU F1 IoU
FC-Siam-Di [108] 86.31 83.31 70.61 54.57
FC-Siam-Conc [108] 83.69 76.77 75.11 60.14
IFNet [109] 88.13 82.93 90.30 71.91
STANet [110] 87.30 77.40 84.12 72.22
BIT [111] 89.31 89.37 88.90 80.01
SNUNet [112] 88.16 87.17 83.89 72.11
ChangeFormer [113] 90.40 88.80 89.83 81.53
RingMo [1] 89.53 82.38 90.90 84.09
RingMo-Aerial 92.71 87.10 94.62 90.09
The comparison results are shown in Tab.X. RingMo-Aerial combined with BIT can achieve change detection results that far exceed other methods. In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection. When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated into our architecture.
The comparison of visual results on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks, as shown in the appendix. The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images. In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
5.2.5 3D-Reconstruction
TABLE XI: 3D-Reconstruction: Comp. on LuoJia-MVS and WHU
@@ 5.2.4 Change Detection
Tab.XI provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets. Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error. On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet. It improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs. 0.141), 0.2% (98.1% vs. 97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a significant improvement. In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%. It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
5.2.6 Object Tracking
TABLE XII: Object Tracking: Comp. on Visdrone-MOT
@@ 5.2.6 Object Tracking
TABLE XIII: Object Tracking: Comp. on AIR-MOT
@@
TABLE XIV: Object Tracking: Comp. on AIR-HSAO
@@
The comparison results on the VisDrone-MOT dataset are shown in Tab.XII. RingMo-Aerial can achieve the best tracking results, far exceeding other methods. In the AIR-MOT dataset (as shown in Tab.XIII), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack. Meanwhile, in the AIR-HSAO dataset (as shown in Tab.XIV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods. Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
Furthermore, the visualization results of the AIR-MOT sequences demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios. The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions. This visualization provides qualitative solid evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in ARS videos.
5.3 Ablation Experiments
5.3.1 Ablation Experiments on pre-training
TABLE XV: Ablation Experiments
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo 71.6 61.9 28.4 45.2
RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
TABLE XVI: Ablation Experiments with the proposed ARS-Adapter
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
We utilize UperNet model on UAVID and Cascade RCNN model on Visdrone-DET. The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL. This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
The comparison is shown in Fig.13 and Tab.XV. The experimental findings reveal a progressive convergence speed and performance enhancement across the three foundation models. Notably, the RingMo-Aerial model trained with MIM and CL exhibits the swiftest convergence and attains the optimal performance. This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pre-training strategy, affirming their mutually reinforcing nature and consistent effectiveness.
In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter. Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL. Comparative results on the UAVID and VisDrone-DET datasets are presented in Tab.XVI. The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
TABLE XVII: Comp. between fine-tuning methods
Methods Parameters UAVID VisDrone-DET
mAcc mIoU mAP mAP50
Froezn 0% 61.2 49.3 21.5 36.2
Full 100% 76.5 66.4 31.1 51.8
Adapter 0.39% 68.9 57.7 28.9 49.1
AdaptFormer 0.39% 71.3 62.1 30.3 50.8
ARS-Adapter * 0.41% 74.3 64.4 31.0 51.4
The experimental results are presented in Tab.XVII. For clarity, we compare ARS-Adapter with full fine-tuning, zero fine-tuning, and two other fine-tuning methods (Adapter and AdaptFormer). Full fine-tuning is a performance upper bound, while zero fine-tuning provides a reference for minimal parameter adjustment. ARS-Adapter requires a similar proportion of parameters(0.41%) to be adjusted as Adapter(0.39%) and AdaptFormer(0.39%), but achieves superior fine-tuning efficiency. Additionally, the performance loss compared to full fine-tuning(100%) is limited to approximately 2%, demonstrating that ARS-Adapter can achieve high performance with minimal parameter adjustments.
6 Conclusion
The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks. Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein. Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks. This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks. In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research. Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment. Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.
Appendix A Semantic Segmentation
A.1 Datasets and Evaluation Metrics
A.2 Experiment Settings
Appendix B Object Detection
B.1 Datasets and Evaluation Metrics
B.2 Experiment Settings
Appendix C Change Detection
C.1 Datasets and Evaluation Metrics
C.2 Experiment Settings
Appendix D Scene Classification
D.1 Datasets and Evaluation Metrics
D.2 Experiment Settings
Appendix E 3D-Reconstruction
E.1 Datasets and Evaluation Metrics
For 3D-Reconstruction, the and WHU [118] datasets are employed.
During the experiment, we used the WHU dataset[118] and the LuoJia-MVS dataset[155] to verify the effectiveness of our proposed method. WHU dataset[118] is the first large-scale aerial multi-view stereo dataset for remote sensing. It was collected from Meitan, Guizhou, China, and was obtained by taking large-scale photos with a tilted five-eye camera on a drone platform. The area of the entire dataset is approximately $6.7\times 2.2km^{2}$, and the image resolution is about 0.1m. In addition, the tilt angle of each camera to the central camera is 40°. The WHU dataset gathers high-rise buildings, mountains, forests, and factory scenes, providing a good benchmark for remote sensing multi-view stereo. LuoJia-MVS[155] dataset is a large-scale multi-view stereo dataset collected from a variety of landform types. It was collected in a part of Baiyun, Guiyang, Guizhou, China. The image size of the entire dataset is $768\times 384$, and the image resolution is about 0.1m. The data consists of five views and pixel-by-pixel depth labels. Land types cover cultivated land, forest land, industrial areas, residential areas, etc.
@@ E.1 Datasets and Evaluation Metrics
E.2 Experiment Settings
Our work is coded based on the PyTorch framework. The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1. During the training process, our method uses the Adam optimizer[156] for parameter optimization and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001. In addition, our method is consistent with the existing multi-stage networks[119]. First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively. Then, deep optimization is performed through three stages of profound hypotheses, and the number of hypotheses is 48, 32, and 8, respectively. During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
Appendix F Object Tracking
F.1 Datasets and Evaluation Metrics
@@ F.2 Experiment Settings
Acknowledgments

### v4

@@ F.1 Datasets and Evaluation Metrics
F.2 Experiment Settings
F.3 Experiment Results
The comparison results are shown in Tab.XVIII. Although RingMo-Aerial does not perform as well as specialized satellite models like SkySense in nadir-view classification, it still demonstrates strong performance, reflecting its robust generalization capability.
TABLE XIX: 3D-Reconstruction: Comp. on LuoJia-MVS and WHU
@@ F.3 Experiment Results
Appendix G 3D-Reconstruction
G.1 Datasets and Evaluation Metrics
For 3D-Reconstruction, the LuoJia-MVS[138] and WHU [136] datasets are employed.
During the experiment, we use three metrics, mean absolute error (MAE), <<0.6m, and <<3-interval, to evaluate the performance of the proposed method.
G.2 Experiment Settings
Our work is coded based on the PyTorch framework. The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1. During the training process, our method uses the Adam optimizer[139] for parameter optimization and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001. In addition, our method is consistent with the existing multi-stage networks[137]. First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively. Then, deep optimization is performed through three stages of profound hypotheses, and the number of hypotheses is 48, 32, and 8, respectively.
G.3 Experiment Results
Tab.XIX provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets. Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error. On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of <<0.6m and <<3-interval accuracy, which is better than the existing advanced method Cas-MVSNet. It improves MAE, <<0.6m, <<3-interval by 13.5% (0.122 vs. 0.141), 0.2% (98.1% vs. 97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a significant improvement. In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a <<0.6m accuracy of 98.0% and a <<3-interval accuracy of 97.9%. It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
Appendix H Object Tracking
H.1 Datasets and Evaluation Metrics
TABLE $X$: Object Tracking: Comp. on Visdrone-MOT
@@ H.1 Datasets and Evaluation Metrics
TABLE XXI: Object Tracking: Comp. on AIR-MOT
@@
TABLE XXII: Object Tracking: Comp. on AIR-HSAO
@@ H.2 Experiment Settings
H.3 Experiment Results
The comparison results on the VisDrone-MOT dataset are shown in Tab.XX. RingMo-Aerial can achieve the best tracking results, far exceeding other methods. In the AIR-MOT dataset (as shown in Tab.XXI), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack. Meanwhile, in the AIR-HSAO dataset (as shown in Tab.XXII), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods. Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
Furthermore, the visualization results of the AIR-MOT sequences (as shown in Fig.20) demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios. The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions. This visualization provides qualitative solid evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in ARS videos.